20230321 UT

Night Info

  • Sunset: 18:35 MST
  • Evening 12-deg twilight: 19:23 MST
  • Morning 12-deg twilight: 05:30 MST
  • Sunrise: 06:17 MST


The entire night was lost due to bad weather.


  • OSA: D. Gonzales-Huerta
  • ISA/SO: A. Cardwell
  • ISp: J. Schuh
  • AO: J.C. Guerra
  • SW: Y. Zhang

Google doc spreadsheet with outline of plan


SW/HW updates

SW patches

  • TCS patch (Yang)
  • AO patch (Xianyu)
  • CCD47 binning patch (Brandon)

HW switch

  • New LUCI Computer (Glenn)

Night Plan

Starting Configuration: LUCI

OSCO (twilight)
LUCI AO (19:30-20:30)
LUCI Bino ESM (20:30-22:30) - 1.5hr
  • Field M67 08:51:45 11:47:46, AOref 10.7, LUCI GS 10.28
  • Script located : /home/lbto/Service_Observing/2023A_D_E/2023A/LUCI/AO_and_ESM_testing/ESM/M67
  • Bino ESM using both LUCI1 and LUCI2. Script generation in ESM_V59.xml OT
  • Scriipt first observes on and off source moving between AO guiding and no AO or active guiding. Then an offset with active guiding is selected towards the end of each filter combination. Use care when returning to source from active guiding as the hotspot for the guider is different from the AOref hotspot and this can put a lot of tip tilt on the mirror and tip tilt guiding may not resume. If only one filter possible, please get BrGamma.
Repoint test (22:30-23:00)

Reconfig LBC

NEO DDT program (23:30-12deg twilight)
  • Scripts: /home/lbto/Service_Observing/2023A_D_E/2023A/LBC/NEO_DDT/March21

Poor weather backup:

Starting Configuration: LBC

LBC/TMS Temperature Test
  • Scripts & Instructions: /home/lbto/Service_Observing/2023A_D_E/2023A/LBC/21Mar_TMS

Night Log

Topic revision: r4 - 21 Mar 2023, AlexanderBecker
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback